LASER INDUCED FLUORESCENCE SPECTROSCOPY OF THE SiNSi ^{2}Δ – X̃ ^{2}Π TRANSITION

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

On our SiCN/SiNC experiment^{a}, we measured laser induced fluorescence excitation spectrum in the range, 27,500 ∼ 33,000 cm^{−1}, under jet cooled condition. Some of the vibronic bands in the spectrum have been assigned to Si containing species, such as Si_{2}, SiCH_{2}, SiCN, and SiNC, but some remain unassigned. The strongest band among them has a ^{2}Π – 2Π type rotational structure. The rotational constants of the ground and excited states are determined to be about 0.11 cm^{−1}, and they are about half of those of SiCN and SiNC. The spin-orbit interaction constant is determined to be 141.7 and 2.7 cm^{−1} for the ground and excited states, respectively. Those are comparable with 140.8 and 4.9 cm^{−1} for the X̃ 2Π and Ã ^{2}Δ states of SiCN ( not SiNC ), respectively. As a result of some discussion, we have finally conclude that the spectrum can be attributed to the ^{2}Δ – X̃ ^{2}Π transition of Si–N–Si. The ^{2}Σ^{+} – X̃ ^{2}Π transition of Si_{2}N was studied by two groups^{b}^{c}. The rotational constant of the ground X̃ ^{2}Π state agrees with that of the previous work. The vibronic structure of the laser induced fluorescence excitation spectrum is unusual; the ^{2}Σ – ^{2}Π type vibronic bands have been only identified.

aM. Fukushima and T. Ishiwata, 60th International Symposium on Molecular Spectroscopy, TC02 (2005).

bD. J. Brugh and M. D. Morse, Chem. Phys. Lett. \underline{267}, 370 (1997).

cZ. J. Jakubek, M. Zachwieja, S. G. Nakhate and B. Simard, 57th International Symposium on Molecular Spectroscopy, MG05